# Are Sounds Sound for Phylogenetic Reconstruction?

### v1

@@ 1 Introduction
ConstenlaChibchan 1214 106 24 0.1 21.71 3.86
CrossAndean 2637 150 19 0.03 28.89 4.32
Dravlex 1341 100 20 0.06 36.85 4.53
FelekeSemitic 2412 150 19 0.05 45.32 4.99
HattoriJaponic 1710 197 10 0.03 34.9 4.47
HouChinese 1816 139 15 0.05 43 6.21
LeeKoreanic 1960 205 14 0.01 36.93 4.31
RobinsonAP 1424 216 13 0.03 24.38 4.51
WalworthPolynesian 6113 207 31 0.05 21.03 4.51
ZhivlovObugrian 1879 110 20 0.04 32.65 3.65
@@ 3.2.1 Bayesian Inference
We initially used an an exponential distribution with a rate of $1.0$ as prior for the $\Gamma$ model of rate heterogeneity. This prior constrains the $\alpha$ shape parameter of the $\Gamma$ distribution to relatively small values. As a consequence, MrBayes obtains $\alpha$ values below 1010 for almost all data sets and posterior samples it draws. This indicates a high to moderate degree of rate heterogeneity. However, our ML analyses (see below) yielded substantially higher ML estimates for $\alpha$ on some datasets. To investigate this discrepancy, we repeated the Bayesian inferences, now using $\mathrm{Uniform}(0.01,100)$ as a prior for the $\Gamma$ distribution of rate heterogeneity. As a consequence, we obtained a different distribution of the $\alpha$ values that better reflects the corresponding ML estimates.
@@ 3.2.2 Maximum Likelihood Tree Inferences
To exclude any potential bias by the selected tree inference method, we also conducted independent Maximum Likelihood (ML) tree inferences. For ML tree inference we used RAxML-NG Kozlov et al. (2019), version 1.2.0. For each dataset and character matrix type (cognate/sound/concatenated) we executed 20 independent ML tree searches using the default tree search configuration of RAxML-NG (10 searches starting from random trees and 10 searches starting from randomized stepwise addition order parsimony trees) under the BIN+G model of binary character substitution with ML estimated base frequencies. We approximate the $\Gamma$ model of rate heterogeneity via four discrete rates. Thus, each inference includes the ML estimate of the $\alpha\in[0.0201,100]$ shape parameter that determines the shape of the $\Gamma$ distribution. The smaller the estimate of $\alpha$, the higher the rate heterogeneity in the respective dataset will be Yang (1995). For three matrices containing cognate data and for two matrices encoding sound correspondences, we obtain an estimate for $\alpha>99.8$, which means that all sites evolve at the same rate and that there is essentially no rate heterogeneity. Hence, trees on these datasets could also be inferred without correcting for rate heterogeneity. For the remaining datasets, the ML estimates of $\alpha$ are below 2020, indicating a moderate to high degree of rate heterogeneity. This extreme bi-modal distribution of $\alpha$ estimates differs substantially from the distribution we observe on tens of thousands of empirical (i.e., non-simulated) molecular datasets Höhler et al. (2021) (also see the respective distribution of alphaalpha values plot at https://github.com/angtft/RAxMLGroveScripts/blob/main/figures/test_ALPHA.png where alphaalpha values range approximately between 0.01 and 1.5).

### v2

@@ 1 Introduction
ConstenlaChibchan (Constenla Umaña, 2005) 1214 106 24 0.1 21.71 3.86
CrossAndean (Blum et al., 2023) 2637 150 19 0.03 28.89 4.32
Dravlex (Kolipakam et al., 2018) 1341 100 20 0.06 36.85 4.53
FelekeSemitic (Feleke, 2021) 2412 150 19 0.05 45.32 4.99
Hattorijaponic (Hattori, 1973) 1710 197 10 0.03 34.9 4.47
HouChinese (Hóu, 2004) 1816 139 15 0.05 43 6.21
LeeKoreanic (Lee, 2015) 1960 205 14 0.01 36.93 4.31
RobinsonAP (Robinson and Holton, 2012) 1424 216 13 0.03 24.38 4.51
WalworthPolynesian (Walworth, 2018) 6113 207 31 0.05 21.03 4.51
ZhivlovObugrian (Zhivlov, 2011) 1879 110 20 0.04 32.65 3.65
@@ 3.2.1 Bayesian Inference
We initially used an an exponential distribution with a rate of $1.0$ as prior for the $\Gamma$ model of rate heterogeneity. This prior constrains the $\alpha$ shape parameter of the $\Gamma$ distribution to relatively small values. As a consequence, MrBayes obtains $\alpha$ values below $10$ for almost all data sets and posterior samples it draws. This indicates a high to moderate degree of rate heterogeneity. However, our ML analyses (see below) yielded substantially higher ML estimates for $\alpha$ on some datasets. To investigate this discrepancy, we repeated the Bayesian inferences, now using $\mathrm{Uniform}(0.01,100)$ as a prior for the $\Gamma$ distribution of rate heterogeneity. As a consequence, we obtained a different distribution of the $\alpha$ values that better reflects the corresponding ML estimates.
@@ 3.2.2 Maximum Likelihood Tree Inferences
To exclude any potential bias by the selected tree inference method, we also conducted independent Maximum Likelihood (ML) tree inferences. For ML tree inference we used RAxML-NG Kozlov et al. (2019), version 1.2.0. For each dataset and character matrix type (cognate/sound/concatenated) we executed 20 independent ML tree searches using the default tree search configuration of RAxML-NG (10 searches starting from random trees and 10 searches starting from randomized stepwise addition order parsimony trees) under the BIN+G model of binary character substitution with ML estimated base frequencies. We approximate the $\Gamma$ model of rate heterogeneity via four discrete rates. Thus, each inference includes the ML estimate of the $\alpha\in[0.0201,100]$ shape parameter that determines the shape of the $\Gamma$ distribution. The smaller the estimate of $\alpha$, the higher the rate heterogeneity in the respective dataset will be Yang (1995). For three matrices containing cognate data and for two matrices encoding sound correspondences, we obtain an estimate for $\alpha>99.8$, which means that all sites evolve at the same rate and that there is essentially no rate heterogeneity. Hence, trees on these datasets could also be inferred without correcting for rate heterogeneity. For the remaining datasets, the ML estimates of $\alpha$ are below $20$, indicating a moderate to high degree of rate heterogeneity. This extreme bi-modal distribution of $\alpha$ estimates differs substantially from the distribution we observe on tens of thousands of empirical (i.e., non-simulated) molecular datasets Höhler et al. (2021) (also see the respective distribution of alphaalpha values plot at https://github.com/angtft/RAxMLGroveScripts/blob/main/figures/test_ALPHA.png where alphaalpha values range approximately between 0.01 and 1.5).